# Complements of caustics of the real $J_{10}$ singularities

### v3

@@ Abstract.
The isotopy classification of Morse perturbations of parabolic real function singularities is completed.
@@ 1. Introduction
This work completes the isotopy classification of Morse perturbations of real parabolic function singularities. The analogous problem for simple singularities was solved in [10], [11], and [14]. For parabolic singularities of classes $X_{9}$ and $P_{8}$, this was done in [15] and [16], respectively. Now, we do the same for the remaining $J_{10}$ singularities. We prove that there are exactly 59 and 56 isotopy classes of Morse perturbations of $J_{10}^{1}$ and $J_{10}^{3}$ singularities, respectively.
@@
In § 2, we describe a combinatorial invariant of isotopy classes of Morse functions that was used in previous works. Then, we compute all admissible values of this invariant for $J_{10}$ singularities and prove that each value can be realized by exactly one or two isotopy classes, depending on whether these classes are invariant or not under a certain symmetry. Finally, we investigate this symmetry condition. Additionally, we realize many isotopy classes by concrete polynomials.
@@
The main invariant of isotopy classes is formulated in the terms of a graph, whose vertices are the collections of certain topological characteristics of particular Morse functions, and whose edges correspond to their standard surgeries. The invariant’s values are the subgraphs of this graph, into which it splits when the edges corresponding to surgeries changing the Morse isotopy class are removed. Our main criterion for the symmetry condition of isotopy classes can be expressed in the terms of one-dimensional cocycles of these subgraphs.
@@
The most interesting part of the classification table obtained is the set of the isotopy classes consisting of polynomials with the maximal possible number of real critical points (ten), and only two different Morse indices. Similar to the $X_{9}$ and $P_{8}$ singularities, almost all of these classes are related to the splittings of the original function singularity into pairs of real critical points with the sum of Milnor numbers equal to ten. The only additional quadruple of isotopy classes (all of which are mapped to each other by certain symmetries) is related in the same way to the canonical extended Dynkin diagram of class $\tilde{E}_{8}$ (which is another name for $J_{10}$ singularities). This situation repeats that of cases $X_{9}$ and $P_{8}$, where the unique exceptional classes are related to the diagrams of types $\tilde{E}_{7}$ and $\tilde{E}_{6}$.
1.2. Main objects and definitions
@@
A function singularity is a germ of a $C^{\infty}$-smooth function at a point where its differential vanishes. A smooth function $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ has a $J_{10}$ singularity at a point $a\in{\mathbb{R}}^{2}$ if in some local coordinates centered at this point it is quasihomogeneous of degree six with weights $\deg x=2$ and $\deg y=1$, and its Milnor number is finite (and then necessarily equal to ten, see [3], [4]). The level set $\{x:f(x)=f(a)\}$ of such a singularity in ${\mathbb{R}}^{2}$ can consist of one or three local branches at $a$. The corresponding subclasses are denoted by $J_{10}^{1}$ and $J_{10}^{3}$.
@@
$J_{10}^{1}$ $\frac{1}{3}(x^{2}+y^{4})(x-\gamma y^{2})$ $\gamma\in(-\infty,\infty)$
$J_{10}^{3}$ $\frac{1}{3}(x^{2}-y^{4})(x-\gamma y^{2})$ $\gamma\in(-1,1)$
The normal forms of $J_{10}$ singularities, to which they can be reduced by a choice of local coordinates, are shown in the last two rows of Table 1. Other rows of this table show the normal forms of simple singularities into some of which the $J_{10}$ singularities can split by small perturbations.
The canonical versal deformations (see [3]) of function singularities of classes $J_{10}^{1}$ and $J_{10}^{3}$ consist respectively of polynomials
(1) $\frac{1}{3}(x-\gamma y^{2})(x^{2}+y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}+3y^{4})y,\qquad\gamma\in(-\infty,+\infty)\ )$
@@
(2) $\frac{1}{3}(x-\gamma y^{2})(x^{2}-y^{4})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad\gamma\in(-1,1),$
@@
The set of non-Morse functions divides these families of polynomials into connected components called isotopy classes of Morse perturbations of $J_{10}$ singularities. Our main goal is the enumeration of these components.
As in [15], [16], we primarily consider slightly greater spaces than the canonical miniversal deformations (1) and (2), which can be reduced to these deformations by an appropriate group of coordinate changes.
Namely, we consider the 16-dimensional space of linear combinations of monomials $x^{\alpha}y^{\beta}$, such that $2\alpha+\beta\leq 6$, with a positive coefficient on the monomial $x^{3}$ and a non-degenerate principal quasihomogeneous part (consisting of monomials with $2\alpha+\beta=6$). This principal part can be of class $J_{10}^{1}$ or $J_{10}^{3}$. The spaces of polynomials with these principal parts are denoted, respectively, $\Phi_{1}$ and $\Phi_{3}$. Again, the sets of Morse functions of classes $\Phi_{1}$ and $\Phi_{3}$ split into connected components (= isotopy classes).
@@
The group $\mathfrak{G}$} consists of diffeomorphisms ${\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ of the form
@@
where $a>0,d>0.$ All coordinate systems in ${\mathbb{R}}^{2}$ obtained from the standard system by the group $\mathfrak{G}$} are called admissible.
The group $\mathfrak{G}$} acts freely on the spaces $\Phi_{1}$ and $\Phi_{3}$ and preserves the set of Morse functions. This group is homotopically trivial (i.e., is diffeomorphic to ${\mathbb{R}}^{6}$), and it reduces all functions of class $\Phi_{1}$ or $\Phi_{3}$ to normal forms (1) or (2).
@@
Each orbit of the action of the group $\mathfrak{G}$} on the space $\Phi_{1}$ ((respectively, $\Phi_{3})$ intersects the space of all polynomials of type ((1)) ((respectively, ((2)))) transversally at a single point.
Proof. The transversality follows from the versality condition of our deformations. Let $f$ be a polynomial of the form (1), and let $\tilde{f}\equiv f\circ G$ for some diffeomorphism $G\in{\mathfrak{G}}$. If the coefficients $a$ and $d$ of this diffeomorphism (see (3)) are not equal to 1 or $c$ is not equal to 0, then the principal quasihomogeneous part of $\tilde{f}$ does not have the standard form of Table 1. If $a=d=1,$ $c=0$, but $\eta\neq 0$ then the coefficient of $\tilde{f}$ at the monomial $xy^{2}$ is non-trivial. If also $\eta=0$ but $b\neq 0$ then the coefficient at $y^{5}$ is not equal to thrice the coefficient at $x^{2}y$. If $\eta=0$ but $\xi\neq 0$, then the coefficient at $x^{2}$ is non-trivial. Thus, $\tilde{f}$ belongs to deformation (1) only if $G=\mbox{Id}$. The proof for $J_{10}^{3}$ is analogous. $\Box$
So, each space $\Phi_{1}$ or $\Phi_{3}$ is canonically diffeomorphic to the direct product of the group $\mathfrak{G}$} and the space of polynomials (1) or (2) ((which is diffeomorphic to ${\mathbb{R}}^{10})$. Additionally, the isotopy classes of Morse functions in spaces $\Phi_{1}$ and $\Phi_{3}$ are in a one-to-one correspondence with the connected components of their intersections with the spaces of polynomials (1) and (2), respectively.
The group $\mathfrak{G}$} also reduces any function of class $\Phi_{1}$ or $\Phi_{3}$ to the normal form with zero coefficients at the monomials $x^{2},x^{2}y,$ and $x^{2}y^{2}$, coefficient 1 at $x^{3}$, and either coefficient $\pm 1$ at $y^{6}$ and coefficient zero at $y^{5}$, or coefficient zero at $y^{6}$ and $xy^{3}$ and coefficient $\pm 1$ at $xy^{4}$.
The main result of the paper is a list of all isotopy classes of Morse polynomials of classes $\Phi_{1}$ and $\Phi_{3}$ or, equivalently, of the form (1) and (2). It is formulated in Theorems 13, 14, and 15.
As a byproduct, we list all possible splittings of real $J_{10}$ singularities into pairs of singular points, such that the sum of their Milnor numbers is ten.
@@
For any pair of simple singularity classes $\Xi$ and $\tilde{\Xi}$ with $\mu(\Xi)+\mu(\tilde{\Xi})=10$, the notation $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{1}$ or $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{3}$ means that there exists a smooth function $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ with a critical point of class $J_{10}^{1}$ or $J_{10}^{3}$ and a one-parametric deformation $F:{\mathbb{R}}^{2}\times[0,\varepsilon)\to{\mathbb{R}}$, $F(\cdot,0)\equiv f,$ of this function, such that, for any $\tau\in(0,\varepsilon)$, the corresponding function $f_{\tau}\equiv F(\cdot,\tau)$ has a critical point of class $\Xi$ and a critical point of class $\tilde{\Xi}$ in such a way that these two critical points depend continuously on $\tau$ and tend to the $J_{10}^{\ast}$ singular point of the function $f$ as $\tau$ tends to 0.
@@ 2. Invariants of isotopy classes of Morse polynomials
This section essentially repeats Section 2 of [15].
@@ 2.1. Trivial invariant
The simplest invariant of isotopy classes of Morse polynomials in ${\mathbb{R}}^{2}$ is their passport, i.e. the triple $(m_{-},m_{\times},m_{+})$ of the numbers of their minima, saddles, and maxima.
By the index considerations, for any Morse polynomial of class $\Phi_{1}$ or $\Phi_{3}$ the sum $M\equiv m_{-}+m_{\times}+m_{+}$ is an even number no greater than ten, and the Euler number $m_{-}-m_{\times}+m_{+}$ is equal to 0 for all Morse polynomials of class $\Phi_{1}$ and to 2-2 for $\Phi_{3}$.
Therefore, when studying a particular class $\Phi_{1}$ or $\Phi_{3}$, we will express the passports by only pairs of numbers, $m_{+}$ and the total number $M$ of real critical points (which determine the remaining passport numbers).

### v5

@@ Abstract.
The complete list of connected components of the set of Morse functions in the deformations of function singularities of class $J_{10}$ is given. Thus, the isotopy classification of Morse perturbations of parabolic real function singularities is finished.
@@ 1. Introduction
This work completes the isotopy classification of Morse perturbations of real parabolic function singularities. The geometry of sets of Morse perturbations of the simplest singularity classes $A_{2}$, $A_{3}$, $A_{4}$ and $D_{4}^{\pm}$ was studied by R. Thom and V. Arnold in the context of catastrophe theoretical problems occurring in biology and optics, see [15], [16], [2], [3]. The local components of the space of Morse functions near all simple function singularities were enumerated in [13], [14], and [19]. For parabolic singularities of classes $X_{9}$ and $P_{8}$, the analogous problem was solved in [20] and [21], respectively. Here, we solve the same problem for the remaining class of parabolic singularities, $J_{10}$. We prove that there are exactly 59 and 56 isotopy classes of Morse perturbations of $J_{10}^{1}$ and $J_{10}^{3}$ singularities, respectively.
@@
In § 2, we recall a combinatorial invariant of isotopy classes of Morse functions (i.e., connected components of the space of these functions) that was used in previous works [19][21]. Then, we compute all possible values of this invariant for $J_{10}$ singularities and prove that each value can be realized by exactly one or two isotopy classes, depending on whether or not these classes are invariant or not under a certain symmetry. Finally, we examine this symmetry condition. Additionally, we realize many isotopy classes by concrete polynomials.
@@
Our main invariant of isotopy classes of Morse functions is formulated in the terms of a graph, whose vertices correspond to the collections of certain topological characteristics of Morse functions, and whose edges correspond to their standard surgeries. The invariant’s values are the subgraphs of this graph, into which it splits when the edges corresponding to those surgeries that can change the Morse isotopy class are removed. Also, the main criterion for the self-symmetry condition of isotopy classes can be expressed in the terms of one-dimensional cocycles of this graph.
@@
The most essential part of the obtained classification table is the list of the isotopy classes of polynomials with the maximal possible number of real critical points (ten), and only two distinct Morse indices. As with the $X_{9}$ and $P_{8}$ singularities, nearly all of these classes are related to the splittings of the original function singularity into pairs of real critical points, the sum of whose Milnor numbers is ten. The only additional quadruple of isotopy classes (all of which are mapped to each other by certain symmetries of the function space) is related in the same way to the canonical extended Coxeter-Dynkin diagram of class $\tilde{E}_{8}$ (another name for $J_{10}$ singularities). This situation repeats that of cases $X_{9}$ and $P_{8}$, where the unique exceptional isotopy classes are related to the diagrams of types $\tilde{E}_{7}$ and $\tilde{E}_{6}$.
1.2. Main objects and definitions (see, e.g., [4])
A point $a\in{\mathbb{R}}^{n}$ is a critical point of a smooth function ${\mathbb{R}}^{n}\to{\mathbb{R}}$ if all first partial derivatives of this function vanish at $a$. A function singularity is a germ of a $C^{\infty}$-smooth function $({\mathbb{R}}^{n},a)\to({\mathbb{R}},0)$ at a point $a\in{\mathbb{R}}^{n}$ where its differential vanishes. Two function singularities at points $a$ and $b$ are equivalent if they can be transformed to each other via a local diffeomorphism $({\mathbb{R}}^{n},a)\to({\mathbb{R}}^{n},b)$ (i.e., they have the same expression in appropriate local coordinate systems centered at $a$ and $b$). The equivalence class of a critical point $(f,a)$ is the equivalence class of the function singularity (ff(a),a).(f-f(a),a). An $l$-parametric deformation of a function singularity $f$ is a family $\{f_{\lambda}\}$ of analytic functions depending on the parameter $\lambda$ running through a neighborhood of the origin point $0\in{\mathbb{R}}^{l}$, such that $f_{0}\equiv f$ and the function $F(x,\lambda)\equiv f_{\lambda}(x)$ in $n+l$ variables $x\in{\mathbb{R}}^{n}$ and $\lambda\in{\mathbb{R}}^{l}$ is regular analytic. The caustic variety of the deformation $F\equiv\{f_{\lambda}\}$ is the set of parameter values $\lambda\in{\mathbb{R}}^{l}$ such that the corresponding function $f_{\lambda}$ has a non-Morse critical point near the origin in ${\mathbb{R}}^{n}$.
@@
A smooth function $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ has a $J_{10}$ singularity at a point $a\in{\mathbb{R}}^{2}$ if in some local coordinates $x$ and $y$ centered at this point it is quasihomogeneous of degree six with weights $\deg x=2$ and $\deg y=1$, and its Milnor number is finite (and then necessarily equal to ten, see [4], [5]). The zero-level set $f^{-1}(0)$ of such a singularity in ${\mathbb{R}}^{2}$ can consist of one or three smooth local branches at $a$. The corresponding subclasses of the $J_{10}$ class are denoted by $J_{10}^{1}$ and $J_{10}^{3}$.
@@
$J_{10}^{1}$ $(x^{2}+y^{4})(x-\gamma y^{2})$ $\gamma\in(-\infty,\infty)$
$J_{10}^{3}$ $(x^{2}-y^{4})(x-\gamma y^{2})$ $\gamma\in(-1,1)$
The normal forms of $J_{10}$ singularities, to which they can be reduced by a choice of local coordinates, are shown in the last two rows of Table 1. Other rows of this table show the normal forms of simple singularities into which the $J_{10}$ singularities can be moved by small perturbations.
The canonical versal deformations (i.e., “sufficiently large” deformations, to which all other deformations can be reduced, see [4]) of function singularities of classes $J_{10}^{1}$ and $J_{10}^{3}$ consist respectively of polynomials
(1) $(x^{2}+y^{4})(x-\gamma y^{2})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}+3y^{4})y,\qquad\gamma\in(-\infty,+\infty),$
@@
(2) $(x^{2}-y^{4})(x-\gamma y^{2})+\lambda_{1}+\lambda_{2}y+\lambda_{3}y^{2}+\lambda_{4}y^{3}+\lambda_{5}y^{4}+\\ \lambda_{6}x+\lambda_{7}xy+\lambda_{8}xy^{2}+\lambda_{9}(x^{2}-3y^{4})y,\qquad\gamma\in(-1,1),$
@@
The caustics in the parameter spaces ${\mathbb{R}}^{9}\times{\mathbb{R}}$ and ${\mathbb{R}}^{9}\times(-1,1)$ of these deformations divide them into connected components called isotopy classes of Morse perturbations of $J_{10}$ singularities. Our main goal is the enumeration of these components.
As in [20], [21], we primarily consider slightly greater spaces than the canonical versal deformations (1) and (2), which can be reduced to these deformations by an appropriate group of diffeomorphisms ${\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$.
Namely, we consider the 16-dimensional space of linear combinations of monomials $x^{\alpha}y^{\beta}$, such that $2\alpha+\beta\leq 6$, with a positive coefficient on the monomial $x^{3}$ and a non-degenerate principal quasihomogeneous part (consisting of monomials $x^{\alpha}y^{\beta}$ with $2\alpha+\beta=6$). This principal part can be of class $J_{10}^{1}$ or $J_{10}^{3}$. The spaces of all polynomials with these principal parts are denoted, respectively, $\Phi_{1}$ and $\Phi_{3}$. Again, the sets of Morse functions of types $\Phi_{1}$ and $\Phi_{3}$ split into connected components (= isotopy classes).
@@
The group $\mathfrak{G}$} consists of all diffeomorphisms ${\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ of the form
@@
where $a>0,d>0.$
This group is diffeomorphic to ${\mathbb{R}}^{6}$. It acts on the spaces $\Phi_{1}$ and $\Phi_{3}$ and preserves the set of Morse functions.
@@
Each orbit of the action of the group $\mathfrak{G}$} on the space $\Phi_{1}$ ((respectively, $\Phi_{3})$ intersects the space of all polynomials of class ((1)) ((respectively, ((2)))) transversally at a single point.
Proof. It is easy to see that any polynomial of class $\Phi_{1}$ or $\Phi_{3}$ can be reduced to the form (1) or (2) by the action of the group $\mathfrak{G}$}. The Lie algebra of this group is generated by the vector fields $x\frac{\partial}{\partial x}$, $y\frac{\partial}{\partial x}$, $y^{2}\frac{\partial}{\partial x}$, $\frac{\partial}{\partial x}$, $y\frac{\partial}{\partial y},$ and $\frac{\partial}{\partial y}$ corresponding to the infinitesimal changes of six parameters of this group. The Lie differentials of polynomials of class (1) or (2) along these vector fields generate the normal bundles of these classes in the spaces $\Phi_{1}$ and $\Phi_{3}$. This implies the transversality of orbits to these classes. Let $f$ be a polynomial of the form (1), and let $G$ be an element of the group $\mathfrak{G}$}. Suppose the polynomial $\tilde{f}\equiv f\circ G$ also has the form (1). Then, the coefficients $a$ and $d$ of the diffeomorphism $G$ are equal to 1 and the coefficient $c$ is equal to 0, since otherwise the principal quasihomogeneous part of $\tilde{f}$ does not have the standard form of Table 1. $\eta=0$ since the coefficient of $\tilde{f}$ at the monomial $xy^{2}$ is trivial. $b=0$ since otherwise the coefficient at $y^{5}$ is not equal to thrice the coefficient at $x^{2}y$. $\xi=0$ since the coefficient at $x^{2}$ is trivial. Thus, $G=\mbox{Id}$. The proof for deformation (2) is analogous. $\Box$
So, each space $\Phi_{1}$ or $\Phi_{3}$ is canonically diffeomorphic to the direct product of the group $\mathfrak{G}$} and the space of polynomials (1) or (2). Additionally, there is a one-to-one correspondence between the isotopy classes of Morse functions in spaces $\Phi_{1}$ and $\Phi_{3}$ and the connected components of their intersections with the spaces of polynomials (1) and (2), respectively.
Alternatively, any function of type $\Phi_{1}$ or $\Phi_{3}$ can be reduced by an element of the group $\mathfrak{G}$} to the normal form with zero coefficients at the monomials $x^{2},x^{2}y,$ and $x^{2}y^{2}$, coefficient 1 at $x^{3}$, and either coefficient $\pm 1$ at $y^{6}$ and zero coefficient at $y^{5}$, or zero coefficient at $y^{6}$ and $xy^{3}$ and coefficient $\pm 1$ at $xy^{4}$.
The main result of the paper is a list of all isotopy classes of Morse polynomials of classes $\Phi_{1}$ and $\Phi_{3}$ or, equivalently, of the form (1) and (2). It is formulated in Theorems 14, 15, and 16.
As a byproduct, we list all possible splittings of real $J_{10}$ singularities into pairs of critical points, such that the sum of their Milnor numbers is ten.
@@
For any pair of simple singularity classes $\Xi$ and $\tilde{\Xi}$ with $\mu(\Xi)+\mu(\tilde{\Xi})=10$, the notation $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{1}$ or $\{\Xi+\tilde{\Xi}\}\rightsquigarrow J_{10}^{3}$ means that there exists a smooth function $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ with a singularity of class $J_{10}^{1}$ or $J_{10}^{3}$ at the origin and a one-parametric deformation $\Theta:{\mathbb{R}}^{2}\times[0,\varepsilon)\to{\mathbb{R}}$, $\Theta(\cdot,0)\equiv f,$ of this function, such that, for any $\tau\in(0,\varepsilon)$, the corresponding function $f_{\tau}\equiv\Theta(\cdot,\tau)$ has a critical point of class $\Xi$ and a critical point of class $\tilde{\Xi}$ in such a way that these two critical points depend continuously on $\tau$ and tend to the origin in ${\mathbb{R}}^{2}$ as $\tau$ tends to 0.
@@
1.3. Lyashko–Looijenga map
The Lyashko–Looijenga map is one of main tools of the proofs and computations in this work. It maps a complexification of the spaces of polynomials (1) and (2) to the space $\mbox{Sym}^{10}({\mathbb{C}}^{1}),$ taking each polynomial to the unordered collection of its critical values. This complexified space consists of all polynomials of the form (2) with arbitrary complex coefficients and unique condition $\gamma\neq\pm 1$. In the restriction to the set of complex polynomials with ten different critical values, this map defines a covering over the configuration space $B({\mathbb{C}}^{1},10)$ of all cardinality 10 subsets in ${\mathbb{C}}^{1}$, see [7], [8].
Both real spaces (1) and (2) can be embedded into this complex space. The embedding of the space (2) is literal. In the case of the space of real polynomials (1), we first identify its parameter space with a real subspace of this complex deformation using the substitution $y=e^{\pi i/4}\tilde{y}$. Namely, this subspace consists of polynomials of the form (2) with coefficients $(\lambda,\gamma)$ such that the numbers $i\gamma,$ $\lambda_{1},$ $e^{\pi i/4}\lambda_{2},$ $i\lambda_{3},$ $e^{3\pi i/4}\lambda_{4},$ $\lambda_{5},$ $\lambda_{6},e^{\pi i/4}\lambda_{7},$ $i\lambda_{8},$ and $\lambda_{9}$ are real. These polynomials are considered as real functions of the real subspace in ${\mathbb{C}}^{2}$ consisting of points $(x,y)$ where $x\in{\mathbb{R}}$, $e^{-\pi i/4}y\in{\mathbb{R}}$.
The restrictions of the Lyashko–Looijenga map to these real spaces send the real polynomials (1) and (2) to the space of point collections in ${\mathbb{C}}^{1}$ that are invariant under the complex conjugation. Over the collections, all of whose points are distinct, these maps are local diffeomorphisms. Also, according to [8], they behave in a proper way near the simplest degenerations of polynomials. This allows us to reduce the topology of spaces of Morse polynomials to the combinatorics of point configurations and associated structures.
@@ 2. Invariants of isotopy classes of Morse polynomials
This section is a shortened version of Section 2 of [20].
@@ 2.1. Trivial invariant
The simplest invariant of isotopy classes of Morse polynomials in ${\mathbb{R}}^{2}$ is their passport, i.e., the triple $(m_{-},m_{\times},m_{+})$ of the numbers of their minima, saddlepoints, and maxima.
According to index considerations, for any Morse polynomial of type $\Phi_{1}$ or $\Phi_{3}$ the sum $M\equiv m_{-}+m_{\times}+m_{+}$ is an even number no greater than ten, and the Euler number $m_{-}-m_{\times}+m_{+}$ is equal to 0 for all Morse polynomials of type $\Phi_{1}$, and to 2-2 for polynomials of type $\Phi_{3}$. Therefore, when studying polynomials of a particular type $\Phi_{1}$ or $\Phi_{3}$, we will express the passports by only pairs of numbers, $m_{+}$ and $M$. These numbers determine the remaining passport numbers.